ThingOne Company has the following information available for the past year. They use machine hours to allocate overhead.

Actual total overhead $70,350
Actual fixed overhead $31,500
Actual machine hours 9,000
Standard hours for the units produced 8,700
Standard variable overhead rate $4.50

What is the variable overhead efficiency variance? Enter the amount as positive number.

Variable overhead efficiency variance $???? Unfavorable
Answer #1

Variable overhead efficiency variance

= (SH - AH) * SR

= (8700 - 9000) * 4.50

= -300 * 4.50

= 1350 Unfavorable

Comment if you face any issues